OpenStudy (anonymous):

A passenger train takes 2 hours less than a freight train takes to travel from Norman to Enid. The passenger train averages 96 km/hr, while the freight train averages 64 km/hr. What is the distance from Norman to Enid?

OpenStudy (anonymous):

@neer2890 @NinjaDevo

OpenStudy (anonymous):

rate x time = distance ( 96 km / h ) T = D ( 64 km / h ) ( T + 2 ) = D

OpenStudy (anonymous):

Setting the equations equal... 96 T = 64 T + 128 32 T = 128 T = 4 hours

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Therefore the distance is 384 km